OpenStudy (anonymous):

Solve the system: x^2 +y^2 =289 2x^2 -y^2 =143 The solutions will have the form (+-a, +- sqrt b) What is b? A. 17 B. 145 C. 3 D. -37 E. none of these

OpenStudy (anonymous):

Do some substitution.

OpenStudy (anonymous):

What should i substitute? can you walk me through this?

OpenStudy (jhannybean):

Just solve this as you would any other system of equations, except this time you will have 2 answers.

OpenStudy (jhannybean):

\[x^2+y^2 = 289\]\[2x^2-y^2=143\]You can see by the process of elimination that the y's can cancel eachother out. You'd be left with:\[3x^2=289+143\]\[3x^2=432\]\[x^2=144\]\[x= \pm 12\] Now we substitute the x values into the first equation: \(x^2+y^2=289\) to find our y value. Here, it does not matter whether you choose to use \(-12\) or \(+12\) because You will end up with a positive value when you substitute it in.

OpenStudy (jhannybean):

So...\[(12)^2 +y^2 = 289\]\[144 +y^2 = 289\]\[y^2 = 289-144 = 145\]\[y = \pm \sqrt{145}\]

OpenStudy (jhannybean):

So now it's evident what your solution can be.

OpenStudy (anonymous):

Okay thanks so much! that makes a lot of sense.